What happens to the Area of a circle if you double the radius? I know the answer is the area would be four fold but I don't know why. I though the area would be doubled.
What happens to the Area of a circle if you square the radius?

Answer by stanbon(75887)

What happens to the Area of a circle if you double the radius? I know the answer is the area would be four fold but I don't know why. I thought the area would be doubled.
---
Let original area = (pi)r^2
---
If you couble the radius, area = (pi)(2r)^2 = 4(pi)r^2
------------------------
What happens to the Area of a circle if you square the radius?
----
If you square the radius, area = (pi)(r^2)^2 = r^2
=============================================
Cheers,
Stan H.